2. Solve using the substitution method. Show your work. If the system has no solution or an infinite number of solutions, state this.
-2x + 0y = -14
x + 11y = 18

please show ne how to do this, not a math genius!

Answer by Alan3354(69443)

-2x + 0y = -14 --> -2x = -14
x = 7
---------
Sub 7 for x
x + 11y = 18
7 + 11y = 18
11y = 11
y = 1